In an experiment with throwing 2 fair dice, consider the events A : The sum of numbers on the faces is 8 B : Doubles are thrown. What is the probability of getting A or B ?
Answer #1

Answer :

Total outcomes for rolling 2 die = 36

Outcomes in event A = {(2,6) (3,5) (4,4) (5,3) (6,2) }

N(A) =5

Outcomes of event B = {(1,1) (2,2) (3,3) (4,4) (5,5) (6,6) }

A union B = { (1,1) (2,2) (3,3) (4,4) (5,5) (6,6) (2,6) (3,5) (5,3) (6,2)}

Probability of A or B = n(A union B)/36 = 10/36 = 5/18